Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to Rs 400 and Rs 120. 


1
Expert's answer
2020-12-07T06:42:25-0500

At equilibrium Qd=Qs

Therefore 1200-P = 120 + 3P

Finding P

1200-120 = 3P+P

1080= 4P

Therefore Equilibriup price = P = Rs 270


At Rs 400, Excess supply = Qs - Qd

Es = 120 + 3(400) - (1200-400)

Es = 1320 - 800 = 500units

Therefore Excess supply = 500units


At Rs 120, Excess demand = Qd - Qs

Ed = (1200-120) - (120 + 3(120))

Ed = 1080 - 480 = 600

Therefore Excess demand = 600units